A departmentstore makes a profit of 1520 on every 25 rolls of film sold. What is the stores profit when 16 rols if film are sold?
Answer by stanbon(75887)

A departmentstore makes a profit of 1520 on every 25 rolls of film sold. What is the stores profit when 16 rols if film are sold?
-----
x/16 = 1520/25
-----
x = 16(60.8)
x = 972.80
================
Cheers,
Stan H.